OpenStudy (anonymous):

The mass of a bucket full of water is 15kg it is being pulled up from 15m deep well. Due to hole in the bucket 6kg water flows out of the bucket . The work done in drawing the bucket out of the well will be

OpenStudy (amistre64):

whats our formula for work?

OpenStudy (anonymous):

plz ans in detail

OpenStudy (amistre64):

no, this site is a study site, not an answer in detail site. your participation is required for a solution to be approached. in other words, you have to do some of the work and thinking.

OpenStudy (anonymous):

i think you also dont know u are just hitting a trial

OpenStudy (amistre64):

so are you asking a valid question, or are you just wasting our time? spamming the site? if its a valid question, then we will need to use the formula for 'work' and theres part of your question that is a bit vague that needs to be cleared up.

OpenStudy (anonymous):

it is a valid question, i mean u shld explain the variable part and upper and lower limit of the intgral

OpenStudy (amistre64):

the variable part is the force portion of the work formula the bucket is losing mass along the way such that the total mass lost is defined. setting up a representative section of it tends to help iron out the details.

OpenStudy (amistre64):

yet its still a good starting point to define the general formulas needed; therefore, what is the formula for work? in general that is.

OpenStudy (anonymous):

\[\int\limits_{?}^{?} F.ds\]

OpenStudy (amistre64):

good, now Force is a measure of weight, so mass * gravity = F mass is changing (15-m) now the vagueness, can we assume the rate of change in mass is constant? 6kg per 15 meters

OpenStudy (anonymous):

lower limit is 0 and upper is 15

OpenStudy (amistre64):

yes, the lower and upper limits are 0 and 15 how dies the mass change with respect to the distance traveled?

OpenStudy (anonymous):

i m not getting that part plz ans

OpenStudy (amistre64):

how much have we lost over 15 meters of travel?

OpenStudy (anonymous):

6kg

OpenStudy (amistre64):

then would you agree that if this is a constant rate of change .. then we can assume m = 6/15 s when s=0, mass loss = 0 when s=15, mass lost is 6 sounds fair?

OpenStudy (amistre64):

\[W=\int_{0}^{15}F~ds\] \[W=\int_{0}^{15}mg~ds\] \[W=\int_{0}^{15}9.8(15-\frac{6}{15}s)~ds\]

OpenStudy (anonymous):

ya i got 1764

OpenStudy (anonymous):

and if g=10 thn ans is 1800 J

OpenStudy (anonymous):

thnx

OpenStudy (amistre64):

good job

OpenStudy (anonymous):

are u a physics lect

OpenStudy (amistre64):

no, im a mathematician. i can do some simple physics calcuations is all :)

OpenStudy (anonymous):

nice